# Perspectives for hyperon and hypernuclei physics

### v1

@@ I Introduction
Nuclei containing strange baryons, so-called Hypernuclei, are unique femto-laboratories for multi-baryon interactions with hyperons Davis (2005); Dalitz (2005); Gal et al. (2016); Chen et al. (2018); Tolos and Fabbietti (2020). Light hypernuclei are particularly interesting, as not only phenomenological models but also ab initio studies based on chiral effective field theories and even lattice quantum chromodynamics calculations are available for such systems Petschauer et al. (2016); Beane et al. (2013).
@@ II.1 The lightest hypernucleus H
Like the deuterium or the tritium nucleus for conventional baryon interactions, a profound understanding of the lightest hypernuclei is a cornerstone for any strong interaction theory dealing with strange baryons Abelev et al. (2010); Chen et al. (2023). A precise and accurate measurement of the hypertriton binding energy gains further relevance from recent hypernuclear studies in ultra-relativistic heavy-ion collisions Adam et al. (2020); Acharya et al. (2023). The large size of the ’ultimate halo nucleus’ H makes it a unique probe for the fragment formation process in ultra-relativistic heavy-ion collisions.
@@ II.3 The case of the neutral nn system
The small binding energy of the hypertriton leads to predictions of the non-existence of bound hypernuclei for isotriplet three-body systems such as nn. However, invariant mass spectroscopy at GSI has reported events that may be interpreted as the bound nn state Rappold et al. (2013). Consistent with the observed lifetime of this system, typical for a weak decay, the measured binding energy is about 1 MeV. Theoretically, this neutral hypernucleus is likely to be unbound, though no firm predictions of its stability can be made. An experimental clarification is required.
@@
Following the tentative observation at GSI, the search for the nn system in the 6Li+12C reaction was repeated with an improved setup at GSI, utilizing the fragment separator facility and the WASA detector. The experiment was successfully performed in 2022. The WASA detector array was originally used at the Svedberg laboratory in Sweden and later at the COSY ring at Forschungszentrum Jülich. Over the past few years, the WASA detector has been installed at the Fragment Separator (FRS) at GSI/FAIR. It was equipped with a new mini draft chamber and additional beam counters. The WASA detector can be opened in the middle to make space for the superconducting solenoid magnet.
The experiment’s goal is to study light hypernuclei during the FAIR Phase 0. The experiment is looking for weak mesonic decays of light hypernuclei produced in 6Li+12C and 12C+12C reactions at 1.96 A GeV. The 12C target is positioned in the mid-focal plane of the fragment separator (FRS). Produced hypernuclei decay weakly behind the primary target. Positive pions are detected in the upgraded WASA detector. The residual nucleus (in the case of the hypertriton 3He) is emitted in the forward direction and is uniquely identified in the second part of the FRS. Thus, the setting of the FRS separator determines the selected hypernucleus.
With a 6Li beam at 1.96 A GeV, three different settings of the fragment separator can be measured, selecting 3He, , or 4He, as well as protons. These settings correspond to the weak pionic decays of H, nn or H, and the free . In addition, with a 12C beam, one more setting can be studied, selecting 3He and simultaneously 9C, corresponding to the decay of H or B. The analysis has started, and the first results are expected soon.
@@ II.4 Study of charge symmetry breaking in hypernuclei
The scattering lengths of proton-proton, proton-neutron, and neutron-neutron were similar in low-energy nuclear experiments, leading to the conclusion that the nuclear force is independent of the charge of the nucleus, called ‘charge symmetry’. The charge symmetry breaking (CSB) was quite small in the interaction. One of the most famous CSB effects was seen in the iso-doublet nuclei 3H and 3He. A 70-keV difference between their binding energies was found after removing the contributions from Coulomb and magnetic effects, which is regarded as the breaking of charge symmetry. This effect can be explained by the mixing effects of and in the interaction.
@@
On the theoretical side, the CSB puzzle also exists. The calculations of the N interaction with mixing can not explain the binding energy difference between H and He in their ground and excited states at the same time Gazda and Gal (2016a). Conversely, this CSB effect can also constrain the scattering lengths of and and their interaction potentials Haidenbauer et al. (2021). Within the four-body cluster model, the structure of the iso-triplet hypernuclei, He, Li, and Be has been studied Hiyama et al. (2009). It was pointed out that a difference of keV between He and Be is expected to originate from the Coulomb force, suggesting that the difference in the hypernuclei system could be small Hiyama et al. (2009). In conclusion, precise measurements of the binding energies in light hypernuclei, particularly H and He, are necessary to understand the CSB mechanism in the N interaction.
@@ II.5 Two- and multi-particle hyperon correlations
Femtoscopy has recently developed into a flourishing field for studying two- and three-body interactions of unbound hadronic systems. Furthermore, this method also provides constraints on the temporal evolution of the emission of various particles (see, e.g., Ref. Vázquez Doce et al. (2025)).
Correlations at small relative momentum measured in p-p and p-Pb collisions at the most violent energy available to this end at the LHC have been exploited to validate for the first time lattice QCD calculations of the nucleon- Acharya et al. (2021), nucleon- Acharya et al. (2019), and nucleon- interactions Acharya et al. (2020a). Very precise data for the p- and p-K correlations have complemented low statistics scattering data with highly precise measurements Acharya et al. (2022, 2020b), and these measurements can be used to constrain chiral effective field theory models more precisely.
The future perspective in this field is to exploit the high statistics expected from the Run 3 campaign at the LHC and the newly upgraded high-rate ALICE detector to investigate the so-far poorly understood nucleon- interaction. The program also includes the investigation of three body forces using the measurement of the momentum correlation for triplets such as p-p- or p-p-, but also exploiting deuteron- and deuteron- correlations to investigate the isospin dependence of the three-body force. This method will thus complement the study of hypernuclei and provide further constraints on theoretical models and input for the equation of state (EoS) of neutron stars, where three-body forces are expected to play a fundamental role.
@@ II.6 Theoretical studies of hyperon-nucleon and hyperon-hyperon interactions
Over the years, many theoretical models have been developed to construct hyperon-nucleon (YN) and hyperon-hyperon (YY) interactions, including various phenomenological models, lattice Quantum Chromodynamics (QCD) simulations, and chiral effective field theories. This section provides a brief review of the development and latest achievements of these three methods.
@@
Similar to the nucleon-nucleon (NN) interaction, the theoretical study of YN and YY interactions began with meson-exchange models, among which the Nijmegen meson-exchange model and the Bonn-Jülich meson-exchange model are the most representative. 1) Nijmegen meson-exchange model. In the 1970s, the Nijmegen collaboration constructed a series of hard-core potentials based on one-boson exchanges (including pseudoscalar mesons, vector mesons, and scalar mesons, etc.) Nagels et al. (1977), achieving simultaneous descriptions of NN and YN scattering. In 1978, Nagels et al. utilized the Regge-pole theory to construct the nucleon-nucleon potential, obtaining nuclear forces based on one-boson exchanges Nagels et al. (1978). In 1989, Maessen et al. extended this to the hyperon-nucleon system using SU(3) flavor symmetry Maessen et al. (1989). Due to its softer behavior near the origin, it was termed the soft-core model (NSC89). In 1997, Rijken et al. considered SU(3) flavor symmetry-breaking effects, introduced mass cutoffs at interaction vertices, and treated the coupling constant ratio as an adjustable parameter, refining and developing the soft-core model and proposing the NSC97 (a-f) models Rijken et al. (1999). These models can accurately describe NN and YN scattering data simultaneously and enable reliable calculations for some -shell hypernuclei. Since then, the Nijmegen collaboration has continuously updated and developed the soft-core model. 2) Bonn-Jülich meson-exchange model. In 1987, the Bonn collaboration proposed the famous Bonn potential model, achieving a high-precision description of NN scattering data. In 1989, Holzenkamp et al. extended the Bonn potential to the YN system, constructing the Jülich89 model Holzenkamp et al. (1989). The subsequently updated Jülich94 version removed the energy-dependent terms in the previous potential Reuber et al. (1994), facilitating convenient and reliable application in the first-order Brueckner calculations of single-particle potentials. In 2005, based on this, Haidenbauer and Meißner constrained the and exchange potentials using a microscopic model of and exchanges and phenomenologically considered short-range contributions from and , proposing the Jülich04 model Haidenbauer and Meissner (2005), which has been also successfully applied in scattering experiments and studies of hypernuclear structures.
@@
Attempts have also been made to describe hyperon-nucleon interactions using quark models. In 1988, the Beijing-Tübingen collaboration proposed a quark cluster model based on one-gluon, pseudoscalar meson, and phenomenological meson exchanges, utilizing the resonating group method to treat 6-quark scattering processes Straub et al. (1988). Based on this method, the modified quark model proposed by Fernández et al. was successfully applied in nuclear force studies Fernandez et al. (1993). Subsequently, the Beijing-Tübingen collaboration extended the modified quark model to the YN system Zhang et al. (1994), simultaneously achieving good descriptions of the NN and YN systems. In 1992, Wang and colleagues proposed the quark delocalization and color screening model by considering fully and incompletely confined configurations in Hilbert space and distinguishing between quark interactions inside and outside hadrons Wang et al. (1992). This model compensated for the lack of intermediate attractive interactions in previous quark models, successfully describing dibaryon interactions and possessing certain predictive capabilities. Another quark model is the Kyoto-Niigata SU(6) quark model proposed by Fujiwara et al. in 1995 Fujiwara et al. (1995). It is based on the full Fermi-Breit interaction, introduces flavor symmetry-breaking effects, and utilizes the resonating group method to treat short-range interactions. Mid- and long-range interactions are modeled using the Nijmegen meson-exchange scheme. Subsequently, the SU(6) quark model has undergone continuous development Fujiwara et al. (2007).
@@
After the 1990s, lattice QCD simulations gradually gained popularity with ever-growing computing power and evolving numerical algorithms. However, due to current computational constraints, lattice QCD can only simulate baryon-baryon interactions at unphysical light-quark masses, where the masses of both mesons and baryons are larger than their physical values. The NPLQCD collaboration and the HAL QCD collaboration are the two main groups studying baryon-baryon interactions using lattice QCD simulations. In 2005, the NPLQCD collaboration proposed using the Lüscher finite-volume method to study YN interactions Beane et al. (2005). In 2009, the HAL QCD collaboration employed the HAL QCD method, where nonlocal potentials are defined within the Nambu-Bethe-Salpeter wave function, to calculate the -wave potential and scattering length for proton- hyperons at MeV Nemura et al. (2009). Since then, these two collaborations have conducted extensive lattice QCD simulations on dibaryon systems in the strangeness sector, but all of these simulations have used a pion mass () larger than 300 MeV. Recently, the HAL QCD collaboration has achieved lattice simulations for dibaryon systems near the physical pion mass ( MeV), such as the system Sasaki et al. (2020).
@@
As a significant milestone in the development of nuclear forces, Weinberg proposed constructing the nucleon-nucleon (NN) interaction using chiral perturbation theory, known as the chiral nuclear force, which has achieved tremendous success. As a natural extension of the NN interaction to the SU(3) flavor space, constructing YN and YY interactions based on chiral effective field theories also holds great potential. In 2005, inspired by the successful study of nuclear forces using the heavy-baryon (HB) chiral effective field theory (ChEFT), the Bonn-Jülich collaboration employed this method to calculate YN and YY scattering observables at leading order (LO) Polinder et al. (2006). The LO potential consists of four-baryon contact terms without derivatives and pseudoscalar meson exchange terms, containing five low-energy constants to be determined. Based on this chiral potential, the scattering amplitude can be obtained by solving the scattering equation nonperturbatively, where renormalization is achieved by introducing an exponential regulator function. The results reproduce the data but with slightly lower precision compared to the phenomenological models such as NSC97f and Jülich04 Rijken et al. (1999); Haidenbauer and Meissner (2005). By further requiring the contact-term potential to satisfy strict SU(3) flavor symmetry, the Bonn-Jülich collaboration constructed leading-order chiral potentials for systems with , , , , and . In 2013, Haidenbauer et al. further constructed YN interactions at the next-to-leading order (NLO) Haidenbauer et al. (2013), achieving a precision comparable to that of established phenomenological models, such as NSC97f and Jülich04. In 2023, Haidenbauer et al. further advanced the YN chiral interaction to the next-to-next-to-leading order (N2LO) by adopting a novel regularization scheme and considering the SU(3) flavor symmetry breaking, which achieved a good description of YN scattering data Haidenbauer et al. (2023). Recently, a study of light hypernuclei was carried out based on the N2LO YN interactions and the no-core shell model Le et al. (2025), where the chiral and three-body forces were treated in a self-consistent way for the first time.

### v2

@@ I Introduction
Nuclei containing strange baryons, so-called Hypernuclei, are unique femto-laboratories for multi-baryon interactions with hyperons Davis (2005); Dalitz (2005); Gal et al. (2016); Chen et al. (2018); Tolos and Fabbietti (2020). Light hypernuclei are particularly interesting, as not only phenomenological models but also ab initio studies based on chiral effective field theories and even lattice quantum chromodynamics calculations are available for such systems Petschauer et al. (2016); Beane et al. (2013); Petschauer et al. (2020).
@@ II.1 The lightest hypernucleus H
Like the deuterium or the tritium nucleus for conventional baryon interactions, a profound understanding of the lightest hypernuclei is a cornerstone of any strong interaction theory dealing with strange baryons Downs and Dalitz (1959); Abelev et al. (2010); Chen et al. (2023). For fitting modern interactions, the H is used to determine the spin dependence Kamada et al. (1998); Sun et al. (2025). A precise and accurate measurement of the hypertriton binding energy gains further relevance from recent hypernuclear studies in ultra-relativistic heavy-ion collisions Adam et al. (2020); Acharya et al. (2023). The large size of the ’ultimate halo nucleus’ H makes it a unique probe for the fragment formation process in ultra-relativistic heavy-ion collisions.
@@ II.3 The case of the neutral nn system
The small binding energy of the hypertriton leads to predictions of the non-existence of bound hypernuclei for isotriplet three-body systems such as nn. However, invariant mass spectroscopy at GSI has reported events that may be interpreted as the bound nn state Rappold et al. (2013a). Consistent with the observed lifetime of this system, typical for a weak decay, the measured binding energy is about 1 MeV. Theoretically, this neutral hypernucleus is likely to be unbound, though no firm predictions of its stability can be made. An experimental clarification is required.
@@
Following tentative observation at GSI Rappold et al. (2013b), the search for the nn system in the 6Li+12C reaction was repeated with an improved setup at GSI, utilizing the fragment separator facility and the WASA detector. The experiment was successfully performed in 2022. The WASA detector array was originally used at the Svedberg laboratory in Sweden and later at the COSY ring at Forschungszentrum Jülich. Over the past few years, the WASA detector has been installed at the Fragment Separator (FRS) at GSI/FAIR. It was equipped with a new mini drift chamber and additional beam counters. The WASA detector can be opened in the middle to make space for the superconducting solenoid magnet. The FRS/WASA is complementary to the time-projection tracker inside the GLAD magnet.
The goal of the experiment is to study light hypernuclei during FAIR Phase 0. The experiment is looking for weak mesonic decays of light hypernuclei produced in 6Li+12C and 12C+12C reactions at a beam energy of 1.96 GeV/nucleon. The analysis is carried out using the invariant mass method Rappold et al. (2013b). The 12C target is positioned in the mid-focal plane of the FRS. Produced hypernuclei decay weakly behind the primary target. Charged pions are detected on the upgraded WASA detector. The residual nucleus (in the case of the hypertriton 3He) is emitted in the forward direction and is uniquely identified in the second part of the FRS. Thus, the setting of the FRS separator determines the selected hypernucleus.
With a 6Li beam at an energy of 1.96 GeV/nucleon, three different settings of the fragment separator can be measured, selecting 3He, , or 4He. These settings correspond to the weak pionic decays of H, nn or H Rappold et al. (2013b). In addition, with a 12C beam, one more setting can be studied, selecting 3He and simultaneously 9C, corresponding to the decay of H or B. The analysis has started, and the first results are expected soon.
@@ II.4 Study of charge symmetry breaking in hypernuclei
The charge symmetry breaking (CSB) was quite small in the interaction Miller et al. (1990). One of the most famous CSB effects was seen in the iso-doublet nuclei 3H and 3He. A 70-keV difference between their binding energies was found after removing the contributions from Coulomb and magnetic effects, which is regarded as the breaking of charge symmetry. This effect can be explained by the mixing effects of and in the interaction.
@@
On the theoretical side, the CSB puzzle also exists Hiyama et al. (2009); Gal (2015); Haidenbauer et al. (2021); Le et al. (2023); Schäfer et al. (2022). The calculations of the N interaction with mixing can not explain the binding energy difference between H and He in their ground and excited states at the same time Gazda and Gal (2016a); Schäfer et al. (2022). Conversely, this CSB effect can also constrain the scattering lengths of and and their interaction potentials Haidenbauer et al. (2021). Within the four-body cluster model, the structure of the iso-triplet hypernuclei, He, Li, and Be has been studied Hiyama et al. (2009). It was pointed out that a difference of keV between He and Be is expected to originate from the Coulomb force, suggesting that the difference in the hypernuclei system could be small Hiyama et al. (2009). Ab-initio calculations within the no-core shell model predicted that the CSB value for the system is small and agrees with the splittings deduced from the empirical binding energies within the experimental uncertainty Le et al. (2023). The computed CSB for the doublet is somewhat larger than the experimental value available Le et al. (2023). In conclusion, precise measurements of the binding energies in light hypernuclei, particularly H and He, are necessary to understand the CSB mechanism in the N interaction.
@@ II.5 Two- and multi-particle hyperon correlations
Femtoscopy has recently developed into a flourishing field for studying two- and three-body interactions of unbound hadronic systems Liu et al. (2025). Furthermore, this method also provides constraints on the temporal evolution of the emission of various particles (see, e.g., Ref. Vázquez Doce et al. (2025)).
Correlations at small relative momentum measured in p-p and p-Pb collisions at the most violent energy available to this end at the LHC have been exploited to validate for the first time lattice QCD calculations of the nucleon- Acharya et al. (2021), nucleon- Acharya et al. (2019a), and nucleon- interactions Acharya et al. (2020a). Very precise data for the p- and p-K correlations have complemented low statistics scattering data with highly precise measurements Acharya et al. (2022, 2020b), and these measurements can be used to constrain chiral effective field theory models more precisely.
The future perspective in this field is to exploit the high statistics expected from the Run 3 campaign at the LHC and the newly upgraded high-rate ALICE detector to investigate the so far poorly understood nucleon- interaction. The program also includes the investigation of the three-particle correlation function using the measurement of the momentum correlation for triplets such as p-p- or p-p-, but also exploiting deuteron- and deuteron- correlations to investigate the isospin dependence of the three-particle correlation function. This method will thus complement the study of hypernuclei and provide further constraints on theoretical models and input for the equation of state (EoS) of neutron stars, where three-body forces are expected to play a fundamental role.
@@ II.6 Theoretical studies of hyperon-nucleon and hyperon-hyperon interactions
Over the years, many theoretical approaches have been developed to construct hyperon-nucleon (YN) and hyperon-hyperon (YY) interactions, including various phenomenological models, lattice Quantum Chromodynamics (QCD) simulations, and chiral effective field theories, which are based on the most general Lagrangian allowed by the assumed symmetries, particularly the (broken) chiral symmetry of QCD. This section provides a brief review of the development and latest achievements of these three methods.
@@
Similar to the nucleon-nucleon (NN) interaction, the theoretical study of YN and YY interactions began with meson-exchange models, among which the Nijmegen meson-exchange model and the Bonn-Jülich meson-exchange model are the most representative. 1) Nijmegen meson-exchange model. In the 1970s, the Nijmegen collaboration constructed a series of hard-core potentials based on one-boson exchanges (including pseudoscalar mesons, vector mesons, and scalar mesons, etc.) Nagels et al. (1977), achieving simultaneous descriptions of NN and YN scattering. In 1978, Nagels et al. utilized the Regge-pole theory to construct the nucleon-nucleon potential, obtaining nuclear forces based on one-boson exchanges Nagels et al. (1978). In 1989, Maessen et al. extended this to the hyperon-nucleon system using SU(3) flavor symmetry Maessen et al. (1989). Due to its softer behavior near the origin, it was termed the soft-core model (NSC89). In 1997, Rijken et al. considered SU(3) flavor symmetry-breaking effects, introduced mass cutoffs at interaction vertices, and treated the coupling constant ratio as an adjustable parameter, refining and developing the soft-core model and proposing the NSC97 (a-f) models Rijken et al. (1999). These models can accurately describe NN and YN scattering data simultaneously and enable reliable calculations for the well depths of 28 MeV in hypernuclei Rijken et al. (1999); Millener et al. (1988); Hasegawa et al. (1996). Since then, the Nijmegen collaboration has continuously updated and developed the soft-core model Nagels et al. (2023). 2) Bonn-Jülich meson-exchange model. In 1987, the Bonn collaboration proposed the famous Bonn potential model, achieving a high-precision description of NN scattering data. In 1989, Holzenkamp et al. extended the Bonn potential to the YN system, constructing the Jülich89 model Holzenkamp et al. (1989). The subsequently updated Jülich94 version removed the energy-dependent terms in the previous potential Reuber et al. (1994), facilitating convenient and reliable application in the first-order Brueckner calculations of single-particle potentials. In 2005, based on this, Haidenbauer and Meißner constrained the and exchange potentials using a microscopic model of and exchanges and phenomenologically considered short-range contributions from and , proposing the Jülich04 model Haidenbauer and Meissner (2005), which has been also successfully applied in scattering experiments and studies of hypernuclear structures.
@@
Attempts have also been made to describe hyperon-nucleon interactions using quark models. In 1988, the Beijing-Tübingen collaboration proposed a quark cluster model based on one-gluon, pseudoscalar meson, and phenomenological meson exchanges, utilizing the resonating group method to treat 6-quark scattering processes Straub et al. (1988). Based on this method, the modified quark model proposed by Fernández et al. was successfully applied in nuclear force studies Fernandez et al. (1993). Subsequently, the Beijing-Tübingen collaboration extended the modified quark model to the YN system Zhang et al. (1994), simultaneously achieving good descriptions of the NN and YN systems. In 1992, Wang and colleagues proposed the quark delocalization and color screening model by considering fully and incompletely confined configurations in Hilbert space and distinguishing between quark interactions inside and outside hadrons Wang et al. (1992). This model addressed the lack of intermediate attractive interactions in previous quark models, successfully describing dibaryon interactions and exhibiting certain predictive capabilities. Another quark model is the Kyoto-Niigata SU(6) quark model proposed by Fujiwara et al. in 1995 Fujiwara et al. (1995). It is based on the full Fermi-Breit interaction, introduces flavor symmetry-breaking effects, and utilizes the resonating group method to treat short-range interactions. Mid- and long-range interactions are modeled using the Nijmegen meson-exchange scheme. Subsequently, the SU(6) quark model has undergone continuous development Fujiwara et al. (2007).
@@
Following the 1990s, lattice QCD simulations gained popularity with the increasing availability of computing power and the development of more advanced numerical algorithms. While significant progress has been made in recent years, high-precision calculations of baryon-baryon interactions at physical pion mass and in large volumes remain computationally challenging. They are an ongoing pursuit in the field. The NPLQCD collaboration and the HAL QCD collaboration are the two main groups studying baryon-baryon interactions using lattice QCD simulations. In 2005, the NPLQCD collaboration proposed using the Lüscher finite-volume method to study YN interactions Beane et al. (2005). In 2009, the HAL QCD collaboration employed the HAL QCD method, where nonlocal potentials are defined within the Nambu-Bethe-Salpeter wave function, to calculate the -wave potential and scattering length for proton- hyperons at MeV Nemura et al. (2009). Since then, these two collaborations have conducted extensive lattice QCD simulations on dibaryon systems in the strangeness sector, but all of these simulations have used a pion mass () larger than 300 MeV. Recently, the HAL QCD collaboration has achieved lattice simulations for dibaryon systems near the physical pion mass ( MeV), such as the system Sasaki et al. (2020).
@@
As a significant milestone in the development of nuclear forces, Weinberg proposed constructing the nucleon-nucleon (NN) interaction using chiral perturbation theory, known as the chiral nuclear force, which has achieved tremendous success. As a natural extension of the NN interaction to the SU(3) flavor space, constructing YN and YY interactions based on chiral effective field theories also holds great potential. In 2005, inspired by the successful study of nuclear forces using the heavy-baryon (HB) chiral effective field theory (ChEFT), the Bonn-Jülich collaboration employed this method to calculate YN and YY scattering observables at leading order (LO) Polinder et al. (2006, 2007). The LO potential consists of four-baryon contact terms without derivatives and pseudoscalar meson exchange terms, containing five low-energy constants to be determined. Based on this chiral potential, the scattering amplitude can be obtained by solving the scattering equation nonperturbatively, where renormalization is achieved by introducing an exponential regulator function. The results reproduce the data but with slightly lower precision compared to the phenomenological models such as NSC97f and Jülich04 Rijken et al. (1999); Haidenbauer and Meissner (2005). By further requiring the contact-term potential to satisfy strict SU(3) flavor symmetry, the Bonn-Jülich collaboration constructed leading-order chiral potentials for systems with , , , , and Polinder et al. (2007); Haidenbauer and Meissner (2010); Haidenbauer et al. (2017). In 2013, Haidenbauer et al. further constructed YN interactions at the next-to-leading order (NLO) Haidenbauer et al. (2013, 2016), achieving a precision comparable to that of established phenomenological models, such as NSC97f and Jülich04. In 2023, Haidenbauer et al. further advanced the YN chiral interaction to the next-to-next-to-leading order (N2LO) by adopting a novel regularization scheme and considering the SU(3) flavor symmetry breaking, which achieved a good description of YN scattering data Haidenbauer et al. (2023). Recently, a study of light hypernuclei was carried out based on the N2LO YN interactions and the no-core shell model Le et al. (2025), where the chiral and three-body forces were treated consistently for the first time.